OpenStudy (anonymous):

How do I solve log(3x + 2) = 3?

OpenStudy (mathstudent55):

Change the log to an exponent.

OpenStudy (australopithecus):

use the rule \[10^{\log_{10}(x)} = x\]

OpenStudy (anonymous):

 10^3 = 1000 (1000 - 2)/3 = 332(2/3)